OpenStudy (anonymous):

integrate Cos(3x)Cos(2x)...any help?

OpenStudy (anonymous):

I think we treat 3x as alpha and 2x as beta, so it's like Integrate Cos(a)cos(b)

OpenStudy (anonymous):

Maybe I need to use some sort of trig formula?

OpenStudy (unklerhaukus):

\[\cos(u)\cos(v)=\frac12\big[\cos(u-v)+\cos(u+v)\big]\]

OpenStudy (anonymous):

hi...can you possibly explain this procedure to me a little?

OpenStudy (anonymous):

How did you get that formula?

OpenStudy (unklerhaukus):

I've just copied one of the Product and Sum Formulas, I can never remember these formulas but I know where to find them .

OpenStudy (anonymous):

It looks like a trig identity but I can't find it in my book or what it would be called

OpenStudy (anonymous):

it's called a sum and product formula?

OpenStudy (unklerhaukus):

what you should do is let 3x=alpha=u 3x=beta=v

OpenStudy (unklerhaukus):

Yeah under trig formulas- sum and product formulas

OpenStudy (anonymous):

so I am doing Integration by Parts on this?

OpenStudy (unklerhaukus):

you dont need to do integration by parts for this problem if you use the tri formula

OpenStudy (anonymous):

oh...I see...I got confused because of the u and v

OpenStudy (anonymous):

quick question...

OpenStudy (anonymous):

is this a sum formula? sin(a+b) = sinacosb+cosasinb

OpenStudy (unklerhaukus):

\[\int \cos(3x)\cos(2x)\mathrm dx\]\[=\frac12\int \cos(3x-2x)+\cos(3x+2x)\mathrm dx\]\[=\]

OpenStudy (unklerhaukus):

i think that \(\sin(a+b) = \sin a\cos b+\cos a\sin b \) is called an angle sum formula

OpenStudy (anonymous):

ok...I think i can integrate it from where you left off...my dumb book totally doesn't have those listed in there though...geesh...but thank you!

OpenStudy (anonymous):

so final answer is 1/2sinx + 1/10sin(5x) +c ?

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]